ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 137, 2003

An, J. and E. A. O’Brien Conjectures on the character degrees of the Harada–Norton simple group HN

157

Anantharaman-Delaroche, C. On spectral characterizations of amenability

1

Ball, K. Entropy and σ-algebra equivalence of certain random walks on random sceneries

35

Baracco, L. Analytic discs in conormal bundles to real submanifolds

of C^{n} 149

Bhattacharya, S. Higher order mixing and rigidity of algebraic actions on compact abelian groups

211

Bhattacharya, S. and K. Schmidt Homoclinic points and isomorphism rigidity of algebraic Z^{d}-actions on zero-dimensional compact abelian groups

189

Bouhjar, K. and J. J. Dijkstra On the structure of n-point sets  321

Bruin, H. and S. van Strien Expansion of derivatives in onedimensional dynamics

223

Bruin, H. and S. Troubetzkoy The Gauss map on a class of interval translation mappings

125

Cooper, S. B., A. Li, A. Sorbi and Y. Yang There exists a maximal 3-c.e. enumeration degree

285

Dijkstra, J. J. See K. Bouhjar

Herzog, M., G. Kaplan and A. Lucchini On subgroups containing non-trivial normal subgroups

183

Ivanov, A. A. Cayley graphs having nice enumerations  61

Kaplan, G. See M. Herzog

Klopsch, B. Enumerating finite groups without abelian composition factors

265

Lee, S. T. and H. Y. Loke Degenerate principal series representations of Sp(p, q)

355

Li, A. See S. B. Cooper

Loke, H. Y. See S. T. Lee

Lucchini, A. See M. Herzog

O’Brien, E. A. See J. An Schmidt, K. See S. Bhattacharya Sorbi, A. See S. B. Cooper Strien, S. van See H. Bruin

Su, Y., K. Zhao and L. Zhu Simple Lie color algebras of Weyl type 109

Troubetzkoy, S. See H. Bruin Yang, Y. See S. B. Cooper Zhao, K. See Y. Su Zhu, L. See Y. Su